OpenStudy (anonymous):

If tan A = 3/4, find the value.(1+cosA)(1-cosA)/(1-sinA)(1+sinA)

OpenStudy (anonymous):

1-cos^2a/1-sin^2a

OpenStudy (anonymous):

what is 1-cos^2a and what is 1-sin^2a ?? this should help

OpenStudy (midhun.madhu1987):

1-cos^2a / 1-sin^2a = ??

OpenStudy (anonymous):

sin^2a/cos^2a=tan^2a=3/4=9/16

OpenStudy (anonymous):

very boring i know this how to solve.but i thought of another method

OpenStudy (anonymous):

see sin(90-theta)=cos,tan(90-theta)=cot,cos(90-theta)=-sin,but i need quadrants 270,180,90,360,,

OpenStudy (anonymous):

sin 23degree without using calculator,

OpenStudy (anonymous):

sin 27degree without using calculator